OpenStudy (anonymous):

Find the vertical asymptotes, if any, of the graph of the rational function. g(x) = x/x^2-16

OpenStudy (anonymous):

\[g(x)=\frac{x}{x^2-16}\] To find the vertical asymptote, we set the denominator equal to zero and solve for x. x^2-16=0

OpenStudy (mathmale):

I'm assuming that your function is\[g(x)=\frac{ x }{ x ^{2}-16 }\] If that's true, follow tomhue's advice, set the den. = to 0, and solve for the 2 values of x that represent the solution of that equation; Write your answers in the form x=a and x=b. The results are the equations of your vertical asymptotes.

OpenStudy (anonymous):

im honestly rearded at this I need the results :/

OpenStudy (anonymous):

retarded

OpenStudy (anonymous):

x^2 -16 = 0 We can use algebra to isolate x. x^2 -16 = 0 x^2 -16 +16 = 0 +16 x^2 = 16 sqrt(x^2) = sqrt(16) Since we're taking a square root, we need to remember the +/- signs. x = +/- sqrt(16) which means x = 4 and x = -4 are the equations of your vertical asymptotes. (x=n is a vertical line, y=n is a horizontal line.)